OpenStudy (anonymous):

Write the slope-intercept form of the equation of the given point withe the given slope. Through: (4, -4), Slope = - 1/3

OpenStudy (cwrw238):

the slope point form is y-y1 = m(x-x1) so plug in m = -1/3 and x1 = 4 and y1 = -4 into this form then rearrange to slope intercept form ( y = mx + c)

OpenStudy (anonymous):

thank you